# Generation of High Order Harmonics in Vacuum for Various Configurations of Interacting Electromagnetic Field

### v1

@@ Abstract
High order harmonic (HOH) generation by interacting extremely intense electromagnetic waves in the quantum vacuum is investigated within the framework of the Heisenberg-Euler formalism. We consider here the process in the lowest order of a perturbation theory relative to the electromagnetic (EM) beam intensity, giving contribution to the HOH generation. The main expressions are obtained for a general geometry, whyle polarizations of different sub-beams forming the EM beam focus are almost the same. Nevertheless, explicit expressions for the HOH generation are derived for the $4\pi$-dipole in-coming waves and for the two crossing Gaussian beams. The former geometry of the EM beam is optimal at a given EM wave power, whereas the latter one is more realistic from the experimental point of view. We consider also a relationship of our present general results with the results, obtained earlier for the HOH generation during of collision of two plane electromagnetic waves.
@@ I Introduction
High power laser developments [1, 2, 3] allow the study of nonlinear physics phenomena related to the electromagnetic field interaction with matter and vacuum [4, 5, 6, 7, 8, 9, 10, 11, 12]. High order harmonics generation in vacuum caused by vacuum polarization effects in the frame of Quantum ElectroDynamics (QED) is one of the more important directions in these investigations [13, 14, 15, 16]. This problem has attracted substantial attention [17, 18, 19, 20, 21, 22, 23, 25, 24, 26, 27, 28] because it discloses the dynamical properties of QED vacuum in strong electromagnetic fields.
There is a number of publications [19, 21, 22, 23, 24, 25, 26, 29] devoted to the detailed theoretical analysis of the harmonic generation. Nevertheless, the theory of this process is far from complete. Several theoretical questions are waiting for their more thorough clarification.
We consider in this paper the HOH generation in the long wave limit for all the waves taking part in the interactions, at least when the wavelengths are larger than the Compton scattering wavelength. Thus, the theoretical consideration will be done within the framework of the Heisenberg–Euler (HE) electrodynamics [30, 31, 32, 33]. We present in this paper a general procedure for calculation of generated HOHs parameters. We take into account a specific form of the Heisenberg–Euler Lagrangian that shifts the HOH generation to higher order of initial laser beam intensity in comparison with naive estimation. The general approach can be applied for a very wide class of incident EM wave geometries.
@@ II Equations of the Heisenberg-Euler electrodynamics
is the Lagrangian of classical electrodynamics with the electromagnetic field tensor $F_{\mu\nu}$ defined in terms of the 4-vector potential $A_{\mu}$ as [41]
@@
$f\left(A_{1},A_{2}\right)$ $=$ $\frac{1}{2}\left(\boldsymbol{B}_{1}\boldsymbol{B}_{2}-\boldsymbol{E}_{1}\boldsymbol{E}_{2}\right)\,;$ (9)
$g\left(A_{1},A_{2}\right)$ $=$ $\frac{1}{2}\left(\boldsymbol{E}_{1}\boldsymbol{B}_{2}+\boldsymbol{E}_{2}\boldsymbol{B}_{1}\right)\,.$ (10)
@@ III.2 Monochromatic in-coming states
$\boldsymbol{A}^{in}=\tilde{\boldsymbol{A}}^{in}+\tilde{\boldsymbol{A}}^{in\,*}\,;\quad\tilde{\boldsymbol{A}}^{in}=e^{-i\omega t}\bar{\boldsymbol{A}}^{in}(\boldsymbol{x})\,\,(\omega>0)\,.$ (33)
@@
where $\Omega=(2n+1)\omega$, and
$\bar{\boldsymbol{j}}_{\Omega}^{out}=\left\langle e^{i\Omega t}\,\boldsymbol{j}^{out}\right\rangle\,.$ (36)
@@
$\boldsymbol{e}_{\boldsymbol{k}}\cdot\bar{\boldsymbol{j}}_{\Omega;\,\boldsymbol{k}}^{out}[\boldsymbol{A}^{in}]=-2\int dV\,\left\langle\left(\frac{\partial\bar{\mathcal{L}}_{1}}{\partial f}\right)\biggr{|}_{\boldsymbol{A}^{in}}f\left(\boldsymbol{A}^{in},\boldsymbol{\mathcal{A}}^{out\,*}\right)+\left(\frac{\partial\bar{\mathcal{L}}_{1}}{\partial g}\right)\biggr{|}_{\boldsymbol{A}^{in}}g\left(\boldsymbol{A}^{in},\boldsymbol{\mathcal{A}}^{out\,*}\right)\right\rangle\,.$ (42)
@@
$H(n,\ell,k)={\color[rgb]{1,0,0}-}\int f^{n+k}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f^{\ell}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,\,f^{k}\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{A}}^{in\,*}\right)f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,$, (45)
$K(n,\ell,k)={\color[rgb]{1,0,0}-}\int f^{n+k+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f^{\ell}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,f^{k}\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{A}}^{in\,*}\right)f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,.$ (46)
@@
${\color[rgb]{1,0,0}-}\int f^{n}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,$, (48)
@@
${\color[rgb]{1,0,0}-}\int f^{n+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,.$ (49)
@@
We recall that this mechanism generates only odd harmonics. We will consider these results in more details below.
We should emphasize that in the case of two perfectly plane waves (of finite duration), considered in Ref. [40], the first non-vanishing effects are related to terms $H(n,n+1,0)$ and $K(n,n,0)$. As a result, intensity of such processes typically are much weaker for a focused laser beams than the processes considered above. More detailed consideration of this interesting question is placed in Sec. V.3.
@@ IV Dipole in-coming EM field
We consider in this section an electric dipole, $E1$, in-coming field. It corresponds to the odd photon state with $j=1$ and $m=j_{z}=0$ [33, 47]. The full set of the photon spherical states $Ej$ and $Mj$ with given quantum numbers $j$ and $m$, corresponding to eigen states of the angular momentum operator, was considered in Refs. [33, 47] and originally in Ref. [48]. A description of this set of the states is reproduced also in App. A.
@@
$E_{\theta}^{in}$ $=$ ${\color[rgb]{1,0,0}-}E_{0}\,e^{-i\omega t}\frac{i}{\omega r}\partial_{r}\left(rj_{1}(\omega r)\right)\sin\theta+\mbox{c.c.}\,$, (52)
@@
where $|E(0)|=4|E_{0}|/3$ is amplitude of oscillation of electric field at $\boldsymbol{x}=0$: $\boldsymbol{E}^{in}(\boldsymbol{x}=0,t)=|E(0)|\,\boldsymbol{e}_{z}\sin(\omega t-\phi)$. We remind that we measure field strengths throughout the whole paper in terms of the critical Schwinger field. To obtain Eq. (59) in physical units we should replace $E(0)$ in it by the ratio $E(0)/E_{S}$.
@@
Figure 2: Dependencies of $\bar{H}_{1}$ (blue line) and $\bar{K}_{1}$ (red line) on the angle $\Theta$. They determine generation of the 3rd harmonics. The lines show real parts of these coefficients. For our normalization of the field $\tilde{\boldsymbol{A}}^{in}$, the imaginary parts vanish.
Figure 3: Dependencies of $\bar{H}_{2}$ (blue line) and $\bar{K}_{2}$ (red line) on the angle $\Theta$. They determine generation of the 5th harmonics. The lines show real parts of these coefficients. For our normalization of the field $\tilde{\boldsymbol{A}}^{in}$ the imaginary parts vanish.
@@ IV.1 Limitations due to electron-positron plasma creation
We need to take into account the process of the electron-positron plasma generation from vacuum, triggered by the Schwinger mechanism [38, 39], and consequently amplified strongly by developing of the $\gamma e^{\pm}$ avalanche [9, 50, 51, 37, 52, 53, 54, 55, 56, 57]. The process will screen the HOH generation considered above. We do not consider here developing of the $\gamma e^{\pm}$ avalanche and formation of the electron-positron plasma triggered by free electrons in the experimental chamber, when the vacuum is not sufficiently perfect.
If $\hbar\omega\sim 0.7$ eV, then probability to create one electron-positron pair in the in-coming wave per one period due to the Schwinger mechanism will be approximately equal to one half at $E(0)/E_{s}\approx 1/14$ [50, 51, 58]. Such in-coming wave carries about $3\times 10^{23}$ photons per period. For higher intensity of the in-coming wave, we will have intense production of the electron-positron plasma [9, 37, 52, 53, 54, 55, 56, 57] that screens the generation of high order harmonics. The expression (59) allows us to estimate the number of quanta of the $(2n+1)$ harmonic generated in this limiting incoming wave per its one period. Rough estimation of this number, $N_{2n+1}(E1)$, is presented in Tab. 2. We conclude that we may hope to observe only the 3rd and 5th harmonics if the intensity of the incoming wave below the threshold of the effective spontaneous creation of the electron-positron plasma from vacuum. Using other polarization of the initial laser beam, we can produce the $M1$ in-coming electromagnetic wave in the focus instead of $E1$. This will mitigate partially the electron-positron plasma production. It would allow us to achieve higher amplitude of the in the focus ($B(0)/E_{s}\simeq 1/10$)) and, hence, higher rate of the HOH generation without significant screening effect, which can be caused by electron-positron plasma production. Calculated analogous values of $N_{2n+1}(M1)$ are also shown in Tab. 2. Nevertheless, this limitation due to electron-positron plasma production remains quite restrictive.

### v2

@@ Abstract
High-order harmonic (HOH) generation during the interaction of extremely intense electromagnetic waves in a quantum vacuum is studied within the Heisenberg-Euler formalism. We consider this process in the first nonvanishing order of perturbation theory. The basic expressions are derived for a general geometry, where the polarizations of the various sub-beams forming the focus of the electromagnetic beam are virtually identical. Explicit expressions for HOH generation are obtained for $4\pi$-dipole incoming wave and for two intersecting Gaussian beams. The former electromagnetic beam geometry is optimal for a given incident electromagnetic wave power, while the latter one is more realistic from an experimental standpoint.
@@ I Introduction
High power laser developments [1, 2, 3] allow the study of nonlinear physics phenomena related to the electromagnetic field interaction with matter and vacuum [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. High order harmonics (HOH) generation in vacuum caused by vacuum polarization effects in the frame of quantum electrodynamics (QED) is one of the more important directions in these investigations [14, 15, 16, 17]. This problem has attracted substantial attention [18, 19, 20, 21, 22, 23, 24, 26, 25, 27, 28, 29] because it discloses the dynamical properties of QED vacuum in strong electromagnetic fields.
There is a number of publications [20, 22, 23, 24, 25, 26, 27, 30] devoted to the detailed theoretical analysis of the harmonic generation. Nevertheless, the theory of this process is far from complete. Several theoretical questions are waiting for their more thorough clarification. For example, the highest possible HOH generation rate in an optimal geometry of the incident electromagnetic (EM) wave remains unclear. One may reveal some lack of explicit analytic expressions for evaluation of the HOH intensity in more or less general but simple geometries.
We consider in this paper the $(2n+1)$-order harmonics generation in vacuum for quasi-monochromatic incident EM wave formally for any $n=1,\,2,\,\dots$. We consider the long wave limit case. It means that the wavelengths of all the waves taking part in the interactions are considerably larger than the Compton scattering wavelength. Thus, the theoretical consideration will be done within the framework of the Heisenberg–Euler (HE) electrodynamics [31, 32, 33, 34]. Lagrangian of the HE electrodynamics contains a perturbation to the classic electrodynamics Lagrangian, and the perturbation can be presented as a sum of terms that can be estimated roughly as $(E/E_{S})^{2m}$ $(m=2,3\dots)$, where $E$ is a typical electric field strength in the the focus of the incident EM wave, and $E_{S}$ is the so called Schwinger critical field, $E_{S}=m_{e}^{2}c^{3}/{e\hbar}=1.32\times 10^{18}$  V/m. The incident wave can be composed experimentally by several laser beams. We present in this paper a general procedure for calculation of generated HOHs parameters. We take into account a specific form of the Heisenberg–Euler Lagrangian that shifts the HOH generation to higher order of initial laser beam intensity in comparison with naive estimation. We show that the intensity of the $(2n+1)$-order harmonic are proportional at least to $(E/E_{S})^{4n+6}$, if we neglect high-orders corrections at $E\ll E_{S}$. For this order in terms of powers of $E/E_{S}$ we take into account only lowest order of the QED small parameter, $\alpha=e^{2}/\hbar c$, the fine structure constant. The general approach can be applied for a very wide class of incident EM wave geometries. Our main goals are the explicit analytic expressions for the HOH intensities for a theoretically optimal geometry of the the incident EM wave, as well as for a geometry that is simplest from both the experimental and theoretical point of view.
@@ II Equations of the Heisenberg-Euler electrodynamics
is the Lagrangian of classical electrodynamics with the electromagnetic field tensor $F_{\mu\nu}$ defined in terms of the 4-vector potential $A^{\mu}=\left(A^{0},\boldsymbol{A}\right)$ as [42]
@@
$f\left(A_{1},A_{2}\right)$ $=$ $\frac{1}{2}\left(\boldsymbol{B}_{1}\cdot\boldsymbol{B}_{2}-\boldsymbol{E}_{1}\cdot\boldsymbol{E}_{2}\right)\,;$ (9)
$g\left(A_{1},A_{2}\right)$ $=$ $\frac{1}{2}\left(\boldsymbol{E}_{1}\cdot\boldsymbol{B}_{2}+\boldsymbol{E}_{2}\cdot\boldsymbol{B}_{1}\right)\,.$ (10)
@@ III.2 Monochromatic in-coming states
$\boldsymbol{A}^{in}=\tilde{\boldsymbol{A}}^{in}+\tilde{\boldsymbol{A}}^{in\,*}\,;\quad\tilde{\boldsymbol{A}}^{in}=e^{-i\omega t}\bar{\boldsymbol{A}}^{in}(\boldsymbol{x})\,\,(\omega>0)\,$, (33)
with $\bar{\boldsymbol{A}}^{in}(\boldsymbol{x})$ obeying in accordance to Eq. (22) the equation:
$\left(\nabla^{2}+\omega^{2}\right)\bar{\boldsymbol{A}}^{in}(\boldsymbol{x})=0\,.$ (34)
@@
where
$\Omega=(2n+1)\omega\,,\mbox{~~~and~~~}\bar{\boldsymbol{j}}_{\Omega}^{out}=\left\langle e^{i\Omega t}\,\boldsymbol{j}^{out}\right\rangle\,.$ (37)
@@
$\boldsymbol{e}_{\boldsymbol{k}}\cdot\bar{\boldsymbol{j}}_{\Omega;\,\boldsymbol{k}}^{out}[\boldsymbol{A}^{in}]=-2\int dV\,\left\langle\left(\frac{\partial\bar{\mathcal{L}}_{1}}{\partial f}\right)\biggr|_{\boldsymbol{A}^{in}}f\left(\boldsymbol{A}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\left(\Omega,\boldsymbol{n}\right)\right)+\left(\frac{\partial\bar{\mathcal{L}}_{1}}{\partial g}\right)\biggr|_{\boldsymbol{A}^{in}}g\left(\boldsymbol{A}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\left(\Omega,\boldsymbol{n}\right)\right)\right\rangle\,.$ (43)
@@
$H(n,\ell,k)=-\int f^{n+k}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f^{\ell}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,\,f^{k}\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{A}}^{in\,*}\right)f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,$, (46)
$K(n,\ell,k)=-\int f^{n+k+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f^{\ell}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,f^{k}\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{A}}^{in\,*}\right)f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,.$ (47)
@@
$\int f^{n}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,$, (49)
@@
$\int f^{n+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}\,.$ (50)
@@
We recall that this mechanism generates only odd harmonics. We will consider these results in more details below. Note, that the result (51) is approximate. The terms containing higher orders of $|F|$, determining the power of the $in$-coming wave, as well as higher orders of $\alpha$ are skipped from this expression. Thus, this expression is exact in this particular sense.
We should emphasize that in the case of two perfectly plane waves (of finite duration), considered in Ref. [41], the first non-vanishing effects are related to the terms $H(n,n+1,0)$ and $K(n,n,0)$. As a result, intensity of such processes typically are weaker for a focused laser beams than the processes considered above. More detailed consideration of this interesting question is placed in Sec. V.3.
It is possible to present $\bar{\boldsymbol{A}}^{in}(\boldsymbol{x})$ in the expressions (49) and (50) as inverse Fourier transform. It would give a momentum representation for $H_{n}$ and $K_{n}$. The momentum representation includes integration over the $(6n+9)$-dimensional 3-momentum space $\left\{\boldsymbol{k}_{1},\dots,\boldsymbol{k}_{2n+3}\right\}$, with a kernel of this integral operator containing several $\delta$-functions: $\prod_{m=1}^{2n+3}\delta\left(\boldsymbol{k}_{m}^{2}-\omega^{2}\right)$ and $\delta\left(\sum_{m=1}^{2n+2}\boldsymbol{k}_{m}-\boldsymbol{k}_{2n+3}-\boldsymbol{k}_{\Omega}\right)$. As a result, this representation is too cumbersome, especially for $n=2,3,\dots$, and is not convenient for our further calculations. We use below only the physical space representation for $H_{n}$ and $K_{n}$ as in Eqs. (49) and (50). Nevertheless, the momentum representation helps to emphasize physical meaning of our expressions. Taking in mind this momentum representation implicitly, we can say that the matrix elements (49) and (50) correspond to the following process expressed in terms of 4-momenta:
$\sum_{m=1}^{2n+1}k_{m}^{\mu}+k_{2n+2}^{\mu}\to k_{2n+3}^{\mu}+k_{\Omega}^{\mu}\,$, (52)
with all the 4-momenta belonging to mass shell of the photons, $k_{\mu}k^{\mu}=0$. The latter condition is automatically fulfilled for the matrix elements $H_{n}$ and $K_{n}$ in the momentum representation due to Eqs. (34) and (42) for $A^{in\,\mu}$ in the Lorentz gauge. Existence in Eq. (52) of the ‘additional’ photons with the momenta $k_{2n+2}^{\mu}$ and $k_{2n+3}^{\mu}$ provides its validity on a non-degenerate $(4n+5)$-dimensional hyper-surface in the $(4n+8)$-dimensional manifold of the photon propagation directions $\left\{\boldsymbol{n}_{1},\dots,\boldsymbol{n}_{2n+3},\boldsymbol{n}_{\Omega}\right\}$ [48]. We consider below wide enough $in$-coming beams in the space of directions of the wave 3-vectors, $\boldsymbol{k}$. For this reason, the mentioned above property of the non-degeneracy provides that the matrix elements (49) and (50) for $n\geq 1$ do not vanish generally. A particular form of this statement corresponding to the case $n=0$ in our notations was previously established in Ref. [23]. This almost evident conclusion will also be discussed below in Sec. V.2 for $n\geq 1$.
A similar consideration but applied to the lower order matrix element $H(n,0,0)$ defined in Eq. (46) leads to its exact vanishing as it is considered above. The latter matrix element corresponds to the process: $\sum_{m=1}^{2n+1}k_{m}^{\mu}\to k_{\Omega}^{\mu}$.
The timelike part of Eq. (52) looks as the following idendity due to definition (37) of $\Omega$:
$\sum_{m=1}^{2n+1}\omega+\omega\to\omega+\Omega\,$, (53)
The first term in the right hand side of this equation, originating from the term $\boldsymbol{k}_{2n+3}$ in Eq. (52), corresponds to the beam of a macroscopic intensity, that is much larger than the intensity corresponding to the unit occupation number. From this point of view, and following Refs. [23, 47], our classic approach treats the emission of the photons $\boldsymbol{k}_{2n+3}$ as a stimulated emission. It will be seen also from our final results for the HOH intensities. The goal of this paper is obtaining of these expressions for the case when the HOH intensities would be high enough, including the case, when typical occupation numbers of quantum states corresponding the beam formed by the vectors $\boldsymbol{k}_{\Omega}$ could be able also to be much higher than 1. For this reason, our classic approach are very reasonable for the perturbation theory for the Lagrangian (1). Applicability of this approach is limited by the condition that the HOH intensity is much less than the intensity of the incident wave, whereas the applicability of the quantum perturbation theory demands formally that typical occupation numbers for the HOHs are much lower than 1. Fortunately, the final results of these two approaches are mainly the same due to the linearity of Eqs. (27) and (36) with respect to $\boldsymbol{A}^{out}$, when their right hand sides are given macroscopic currents.
@@ IV Dipole in-coming EM field
We consider in this section an electric dipole, $E1$, in-coming field. It corresponds to the odd photon state with $j=1$ and $m=j_{z}=0$ [34, 49]. Here, $j$ is the total angular momentum of the photon from this EM wave, and $m$ is its $z$-component. The full set of the photon spherical states $Ej$ and $Mj$ with given quantum numbers $j$ and $m$, corresponding to eigen states of the angular momentum operator, was considered in Refs. [34, 49] and originally in Ref. [50]. A description of this set of the states is reproduced also in App. A. The set was constructed in Refs. [34, 49, 50] from the following representation of the exact solutions of the Maxwell equations (in the transverse gauge):
$\boldsymbol{A}=\frac{e^{-i\omega t}}{(2\pi)^{2}}\int\delta(|\boldsymbol{k}|-\omega)\,\boldsymbol{a}_{\boldsymbol{k}}\,e^{i\boldsymbol{k}\boldsymbol{x}}\,d\boldsymbol{k}+\mbox{c.c.}\,.$ (54)
Here, $\omega$ is frequency of the monochromatic wave, and $\boldsymbol{a}_{\boldsymbol{k}}$ is tangent to the sphere $|\boldsymbol{k}|=\omega$ in the $\boldsymbol{k}$-space.
@@
$E_{\theta}^{in}$ $=$ $-E_{0}\,e^{-i\omega t}\frac{i}{\omega r}\partial_{r}\left(rj_{1}(\omega r)\right)\sin\theta+\mbox{c.c.}\,$, (56)
@@
where $|E(0)|=4|E_{0}|/3$ is amplitude of oscillation of electric field at $\boldsymbol{x}=0$: $\boldsymbol{E}^{in}(\boldsymbol{x}=0,t)=|E(0)|\,\boldsymbol{e}_{z}\sin(\omega t-\phi)$. Note, that the expression (65) does not contain any parameters, besides $n$. We remind that we measure field strengths throughout the whole paper in terms of the critical Schwinger field. To obtain Eq. (63) in physical units we should replace $E(0)$ in it by the ratio $E(0)/E_{S}$.
@@
Figure 2: Dependencies of $\bar{H}_{1}$ (blue) and $\bar{K}_{1}$ (red) on the angle $\Theta$. They determine generation of the 3rd harmonics. The lines show real parts of these coefficients. For our normalization of the field $\tilde{\boldsymbol{A}}^{in}$, the imaginary parts vanish.
Figure 3: Dependencies of $\bar{H}_{2}$ (blue) and $\bar{K}_{2}$ (red) on the angle $\Theta$. They determine generation of the 5th harmonics. The lines show real parts of these coefficients. For our normalization of the field $\tilde{\boldsymbol{A}}^{in}$ the imaginary parts vanish.
@@ IV.1 Limitations due to electron-positron plasma creation
We need to take into account the process of the electron-positron plasma generation from vacuum, triggered by the Schwinger mechanism [39, 40], and consequently amplified strongly by developing of the $\gamma e^{\pm}$ avalanche [10, 52, 53, 38, 54, 55, 56, 57, 58, 59, 60]. The process will screen the HOH generation considered above. We do not consider here developing of the $\gamma e^{\pm}$ avalanche and formation of the electron-positron plasma triggered by free electrons in the experimental chamber, when the vacuum is not sufficiently perfect.
If $\hbar\omega\sim 0.7$ eV, then probability to create one electron-positron pair in the in-coming wave per one period due to the Schwinger mechanism will be approximately equal to one half at $E(0)/E_{s}\approx 1/14$ [52, 53, 61]. Such in-coming wave carries about $3\times 10^{23}$ photons per period. For higher intensity of the in-coming wave, we will have intense production of the electron-positron plasma [10, 38, 54, 55, 56, 57, 58, 59, 60] that screens the generation of high order harmonics. The expression (63) allows us to estimate the number of quanta of the $(2n+1)$ harmonic generated in this limiting incoming wave per its one period. Rough estimation of this number, $N_{2n+1}(E1)$, is presented in Tab. 2. We conclude that we may hope to observe only the 3rd and 5th harmonics if the intensity of the incoming wave below the threshold of the effective spontaneous creation of the electron-positron plasma from vacuum. Using other polarization of the initial laser beam, we can produce the $M1$ in-coming electromagnetic wave in the focus instead of $E1$. This will mitigate partially the electron-positron plasma production. It would allow us to achieve higher amplitude of the in the focus ($B(0)/E_{s}\simeq 1/10$)) and, hence, higher rate of the HOH generation without significant screening effect, which can be caused by electron-positron plasma production. Calculated analogous values of $N_{2n+1}(M1)$ are also shown in Tab. 2. Nevertheless, this limitation due to electron-positron plasma production remains quite restrictive.